THE ELECTRIC DIPOLE MOMENT OF IRIDIUM MONOSILICIDE, IrSi

ANH LE AND \underline{TIMOTHY C. STEIMLE}, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; LAN CHENG AND JOHN F. STANTON, The University of Texas at Austin, Austin, TX 78712-0165.

The optical spectrum of iridium monosilicide (IrSi) was recently observed using REMPI spectroscopy in the range 17200 to 23850 cm^{−1}^{a}. The observation was supported by an ab initio calculation which predicted a X^{2}Δ_{5/2} state. Here, we report on the analysis of the optical Stark effect for the X^{2}Δ_{5/2} and [16.0]1.5 (v=6) states. The (6,0)[16.0]1.5 - X^{2}Δ_{5/2} and the (7,0)[16.0]3.5- X^{2}Δ_{5/2} bands of IrSi have been recorded using high-resolution laser-induced fluorescence spectroscopy. The observed optical Stark shifts for the ^{193}IrSi and ^{191}IrSi isotopologues were analyzed to produce the electric dipole moments of -0.4139(64)D and 0.7821(63)D for the X^{2}Δ_{5/2} and [16.0]1.5 (v=6) states, respectively. The negative sign of electric dipole moment of the X^{2}Δ_{5/2} state is supported by high-level quantum-chemical calculations employing all-electron scalar-relativistic CCSD(T) method augmented with spin-orbit corrections as well as corrections due to full triple excitations. In particular, electron-correlation effects have been shown to be essential in the prediction of the negative sign of the dipole moment. A comparison with other iridium containing molecules will be made.

aMaria A. Garcia, Carolin Vietz, Fernando Ruiprez, Michael D. Morse, and Ivan Infante, Kimika Fakultatea, Euskal Herriko. J.

Chem. Phys., (submitted)